OpenStudy (anonymous):

The formula S= 2 pi rh+ 2 pi r^2 gives the surface area S of a cylinder with height h and radius r. Solve the formula for h. Find h if r=5 cm and S=400 cm square.

OpenStudy (aum):

\(S= 2 \pi rh+ 2 \pi r^2\) To solve for h, isolate h Subtract \(2\pi r^2\) from both sides: \(S - 2\pi r^2 = 2\pi rh\) Divide both sides by \(2\pi r\) \(\Large \frac{S - 2\pi r^2}{2\pi r} = h\)

OpenStudy (anonymous):

Here you go!

OpenStudy (anonymous):

thanks so much guys :D ur a great help!!

OpenStudy (aum):

Before plugging in the numbers they are asking you to solve for h first. So do the steps in my first reply. Then follow the image posted by phuongduongm.

OpenStudy (anonymous):

ohhhh okok i get it now. you guys are so helpful thanks so much :)